OpenStudy (anonymous):

Can anyone help me with the rule of descartes and the rational root theorem?

OpenStudy (anonymous):

Please!!!?

OpenStudy (anonymous):

is there a specific question?

OpenStudy (anonymous):

not really but these theorems are really hard to grasp can you help?

OpenStudy (anonymous):

it is much easier to explain by example, but if you want a very easy to understand (color coded) explanation of descartes rule of sign, with examples, try here http://www.purplemath.com/modules/drofsign.htm

OpenStudy (anonymous):

Thank you! But, what about the rational root theorem?

OpenStudy (anonymous):

purplemath also has an excellent explanation of that as well

OpenStudy (anonymous):

although that one is easy enough the possible rational roots of a polynomial are numbers \(\frac{p}{q}\) where \(p\) divides the constant and \(q\) divides the leading coefficient so for example if it is \[f(x)=2x^4+ stuff +5\] then the possible rational roots are \[\pm1,\pm5,\pm\frac{1}{2},\pm\frac{5}{2}\]

OpenStudy (anonymous):

Thank you so much I understand now :)